OpenStudy (anonymous):

The graph shows the amount of time that Dewayne spent on a practice exam based on the number of questions he completed. What is the function rule? How much time will he spend if the actual exam has 20 questions?

OpenStudy (anonymous):

I little hard to see

OpenStudy (anonymous):

100 minutes

OpenStudy (anonymous):

take those points, find the slope. so between (2,10) and (4,20) slope is 5. then put the slope in that (y-y1)=m(x-x1) formula and get the general rule y=5x (I think...)

OpenStudy (anonymous):

and yea, it ends up being 100 mins ;)

OpenStudy (anonymous):

y = x/2 or y = 5x

OpenStudy (anonymous):

y = 5x

OpenStudy (anonymous):

y = 5x probably